OpenStudy (anonymous):

Quick Integral Question: (cos(x))^5 / sin(x) The cos^5(x)...the x is not raised to the 5th the cosine is.... How would you find the integral of that. What rules should I use?

OpenStudy (dumbcow):

substitution u = cos x du = -sin x dx giving \[\int\limits \frac{u^5}{u^2 - 1} du\]

zepdrix (zepdrix):

But the sine is in the denominator cow 0_o I'm confused how you got that.

OpenStudy (dumbcow):

from here use long division \[\rightarrow \int\limits u^3 + u + \frac{u}{u^2 - 1}\] finally split up last fraction using partial fractions \[\frac{u}{u^2 -1} = \frac{u}{(u+1)(u-1)} = \frac{A}{u+1} + \frac{B}{u-1}\]

OpenStudy (dumbcow):

@zepdrix, dx = du/-sin yielding "-sin^2" in denominator which equals "cos^2 - 1"

zepdrix (zepdrix):

Oh ok :) Interesting. Here is how I would approach it personally. Just another idea. Break the cos^5 into cos^4 and cos, write the cos^4 as cos^2 being squared, \[\Large\rm \int\limits \frac{(\cos^2 x)^2 \cos x ~dx}{\sin x}\]Then apply your Pythagorean Identity:\[\Large\rm \int\limits\limits \frac{(1-\sin^2x)^2 \cos x ~dx}{\sin x}\] Then let \(\Large\rm u=\sin x, \qquad du=\cos x~dx\)

OpenStudy (anonymous):

so the denominator and the cosx cancel out leaving me with u^2?

OpenStudy (dumbcow):

@JohnathonNY , both methods will work but i would use zepdrix here, simpler steps haha i over complicated it :)

OpenStudy (dumbcow):

no it would leave you with \[\int\limits \frac{(1-u^2)^2}{u} du\] expand, then divide each term by "u"